Bob is 10 years older than kim. The sum of their age is 60. How old is bob
Answer by ikleyn(52795)

.
b - k = 10,
b + k = 60.

Add the two equations. You will get

2b = 10+60 = 70.

Hence, b = 35. <<<===Bob's age.